Suppose Charlie O'Brian of the Braves hits a baseball straight upward at 150ft/sec from a height of 5 ft.
If an object is given an initial velocity of v feet/second from a height of s feet, then its height S after t seconds is given by the formula
S= -16t^2 + vt + s.
a. Use the formula to determine how long it takes the ball to return to earth.
b. Estimate the maximum height reached by the ball.

Answer by stanbon(75887)

Suppose Charlie O'Brian of the Braves hits a baseball straight upward at 150ft/sec from a height of 5 ft.
If an object is given an initial velocity of v feet/second from a height of s feet, then its height S after t seconds is given by the formula
S(t) = -16t^2 + vt + s.
a. Use the formula to determine how long it takes the ball to return to earth.
At earth its height will be zero.
-16t^2 + 150t + 5 = 0
t = [-150 +- sqrt(150^2 - 4*-16*5)]/-32
t = [-150 +- sqrt(22820)]/-32
Positive solution:
t = [-150 - 151.063]/-32
t = 9.408 seconds
-----------------------------
b. Estimate the maximum height reached by the ball.
Max occurs when t = -b/2a = -150/[-32] = 4.6875 seconds
S(4.6875) = -16(4.6875)^2 + 150(4.6875) + 5 = 356.56 ft.
==============================================================
Cheers,
Stan H.